3.) A gas takes up 14.8 liters of 24°C. What temperature in kelvin is required to obtain a volume of 25.0 liters at constant pressure?

4.) How many moles of chlorine gas are present at 25°C, 0.9 atm, with a volume of 14.2 L? S.)

If "Bor" has a sample of gas that has a volume of 8.2 1liter at 25°C and 2 atm, how much volume will it take up if you decrease pressure to 1.5 atms and increase temperature to 100°C?

6.) 25 liters of gas A is pumped into a container at 25°C and 760 torr with 20 liters of gas B at 25°C and 700 torr. Calculate the total pressure when both gases are pumped into a tank with 10 liters at 25°C

Please
Answer #1

3) Given that she 14.8L 2.7 25.00 T-24c =297K T2 = ? According to Charles low TT 14.8 28.OL . 292 TL ar.oLX 2941 1 14.8 L 01

V=2194 13.7L b . Asmie present at 25°c, calm helh 20t 25 Lip 2 6) 20L y = zich T-25% = 298k. P = (alm. pv (atmx 25 No. of mol

Final volume = 13.7 L

6)

Given that

Temperature of gas A = 25°C = 298 K

Pressure of gas A = 760 torr = 1 atm

Volume of gas A = 25 L

Number of moles of A = PV/RT

= 1 atm * 25 L /(0.0821 L atm /mol K)*298 K

= 1.02 moles

Given that

Temperature of gas B = 25°C = 298 K

Pressure of gas B = 700 torr = 0.921 atm

Volume of gas B = 20 L

Number of moles of A = PV/RT

= 0.921 atm * 25 L /(0.0821 L atm /mol K)*298 K

= 0.75 moles

Total number of moles of A &B = 1.77 moles

Gaseous mixture is present at 298 K and occupies a volume of 10 L

Pressure of the mixture = nRT/V

= 1.77 moles *0.0821 L atm/mol K/10 L

= 4.33 atm